# Improving Discrete Diffusion Unmasking Policies Beyond Explicit Reference Policies

### v1

@@ Abstract
Masked diffusion models (MDMs) have recently emerged as a novel framework for language modeling. MDMs generate sentences by iteratively denoising masked sequences, filling in [MASK] tokens step by step. Although MDMs support any-order sampling, performance is highly sensitive to the choice of which position to unmask next. Prior work typically relies on rule-based schedules (e.g., max-confidence, max-margin), which provide ad hoc improvements. In contrast, we replace these heuristics with a learned scheduler. Specifically, we cast denoising as a KL–regularized Markov decision process (MDP) with an explicit reference policy and optimize a regularized objective that admits policy-improvement and convergence guarantees under standard assumptions. We prove that the optimized policy under this framework generates samples that more closely match the data distribution than heuristic schedules. Empirically, across four benchmarks, our learned policy consistently outperforms max-confidence: for example, on Sudoku, where unmasking order is critical, it yields a 20.1% gain over random and a 11.2% gain over max-confidence.
@@ 1 Introduction
Related works. Reinforcement learning (RL) has emerged as an effective tool to fine-tune large language models toward complex objectives or preferences (Ouyang et al., 2022; Guo et al., 2025; Shao et al., 2024; Rafailov et al., 2024). In MDMs, Zekri & Boullé (2025) apply policy gradient methods to a pretrained discrete diffusion model so as to maximize a task-specific reward at the last denoising step. Zhao et al. (2025) adopts the GRPO framework to improve the reasoning ability by rewarding only the final answer accuracy. Zhu et al. (2025) propose a variance reduced preference optimization (VRPO) for MDMs, which shares the concept of direct preference optimization (DPO) (Rafailov et al., 2024) in ARMs. While the aforementioned works utilize RL to post-train the MDM itself, they differ from ours, where we train the unmasking policy model. Recently, Huang et al. (2025) introduce DColT, which also trains an unmasking policy via reinforcement learning. Our approach differs in that we formulate denoising as a KL–regularized MDP with an explicit reference policy and optimize a regularized objective that admits policy-improvement and convergence guarantees under standard assumptions. Furthermore, while Huang et al. (2025) jointly applies RL to train both the unmasking policy and the underlying MDMs for maximum accuracy, our work focuses solely on optimizing the unmasking policy for frozen MDMs, thereby reducing computational cost and mitigating the risk of reward over-optimization.
@@ 2 Preliminary
Notations. Let the distribution $p_{\rm{data}}$ be the data distribution over sequences of length $L$ with vocabulary $\mathcal{V}=\{1,...,m\}$, and let the set of all possible sequences be $\mathcal{X}=\mathcal{V}^{L}$. We define $p_{L}$ as the Kronecker delta distribution that assigns all probability mass to the sentence in which every token is masked. We denote the kernel forms of $p_{\rm{data}}$ and $p_{L}$ by ${\mathbf{P}}_{\rm{data}}$ and ${\mathbf{P}}_{L}$, respectively, with entries ${\mathbf{P}}_{\rm{data}}]{{\bm{x}}}=p_{\rm{data}}({\bm{x}})$. We write ${\bm{x}}_{n}=\{x_{n}^{1},x_{n}^{2},\dots,x_{n}^{L}\}$ for the $n$-th step sequence with $L$ tokens. We denote the mask by $\mathbb{M}$ and the set of sequences with $n$ masks by $\mathcal{X}_{n}=\{{\bm{x}}|{\bm{x}}\in\{\mathcal{V}\cup\{\mathbb{M}\}\}^{L},\sum_{i=1}^{L}{\textbf{1}(x^{i}=\mathbb{M})}=n\}$, so that ${\bm{x}}_{n}\in\mathcal{X}_{n}$ has $n$ masks and we refer $a^{i}[{\bm{x}}]$ for the $i$-th mask index of $\bm{x}$}. For example, if the number of masks in $\bm{x}$} is $n$, then $a^{i}[{\bm{x}}]\in\{1,\dots,L\}$, $x^{a^{i}[{\bm{x}}]}=\mathbb{M}$ and $i\in\{1,\dots,n\}$. For brevity, we will omit $\bm{x}$} in $a^{i}[{\bm{x}}]$ for the rest of the paper. We denote MDMs as $\theta$ and their model distribution as $\pi_{\theta}$. We denote Markov transition kernel ${\mathbf{T}}:\mathcal{X}\times\mathcal{X}\to[0,1]$ where ${\mathbf{T}}({\bm{x}},{\bm{x}}^{\prime})$ is the probability of $\bm{x}$} transform into $\bm{x}$$^{\prime}$ such that $\sum{\mathbf{T}}_{{\bm{x}}^{\prime}\in\mathcal{X}}({\bm{x}},{\bm{x}}^{\prime})=1$ for all $\bm{x}$}.
@@
Now consider the MDM directly answer the questions, filling the mask of sentence: “House 1: name: [MASK], food: [MASK], House 2: name: [MASK], food: [MASK]”. Obviously, if MDM tries to answer the name of the person who lives in the first house, the problem becomes trivial. However, if it tries to fill attributes for the second house first before filling attributes of the first house, the problem becomes much harder. As the number of questions and clues grows, this kind of discrepancy would be much larger. Indeed, the zebra puzzle is also known as an NP-hard problem (Shah et al., 2024; Yato & Seta, 2003). Note that this is a highly simplified example to help understand the problem of MDM training, and refer to Kim et al. (2025) for a complete explanation.
@@ 3 Improving Unmasking Policy Beyond a Reference Policy
3.1 Existence of optimal path Beyond $g_{\rm{conf}}$
@@
We test various $g$ in GSM8K (Cobbe et al., 2021), the mathematical questions benchmark, to probe whether max-confidence is merely strong or whether substantially better unmasking paths exist. We compare three schedulers: $g_{\rm conf}$, $g_{\rm rand}$, and $g_{\rm{Top-K}}$. Here, $g_{\rm conf}$ is deterministic, while $g_{\rm Top\text{-}K}$ smoothly interpolates between random and max-confidence.
@@ 3.3 Theoretical analysis
Assume there exists ideal MDM $\theta^{*}$ satisfying $\mathcal{L}_{\rm{MDM}}(\theta^{*})=0$ for $p_{\rm{data}}$ where data is composed of prompt and oracle answer $\{{\mathbf{q}},{\mathbf{a}}\}$ such that $r({\mathbf{q}},{\mathbf{a}})=1$. Now consider incomplete MDM $\theta$ such that $\mathcal{L}_{\rm{MDM}}(\theta)>0$, and $g_{\phi^{*}}$ that satisfies Eq. (15) for $\theta$. Let let $g_{\mathrm{ref}}$ be a policy model such that, for every $\mathbf{q}$}, $p_{g_{\rm{ref}}}({\bm{x}}_{0}\mid{\mathbf{q}})>0$ whenever $p_{\theta^{*}}({\bm{x}}_{0}\mid{\mathbf{q}})>0$, i.e., $\operatorname{supp}(p_{\theta^{*}}(\cdot\mid{\mathbf{q}}))\subseteq\operatorname{supp}(p_{g_{\mathrm{ref}}}(\cdot\mid{\mathbf{q}}))$. Then, the following inequality holds:
$D_{\rm{KL}}(p_{\rm{\theta^{*}}}({\bm{x}}_{0}|{\mathbf{q}})||p_{g_{\phi^{*}}}({\bm{x}}_{0}|{\mathbf{q}}))\leq D_{\rm{KL}}(p_{\rm{\theta^{*}}}({\bm{x}}_{0}|{\mathbf{q}})||p_{g_{\rm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})).$ (16)
@@
$D_{\rm{KL}}(p_{\rm{data}}({\mathbf{a}}|{\mathbf{q}})||p_{g_{\phi^{*}}}({\bm{x}}_{0}|{\mathbf{q}}))\leq D_{\rm{KL}}(p_{\rm{data}}({\mathbf{a}}|{\mathbf{q}})||p_{g_{\rm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})).$ (17)
@@
Mroueh (2025) provides policy dynamic under training Eq. (12). Using the result given in Lemma 1 in Appendix, we can express $g_{\phi^{*}}$ with $g_{\rm{ref}}$ and thereby derive KL divergence. Please refer to Appendix C.1 for detailed proof. ∎
@@ 3.4 Practical realization of unmasking policy optimization
Although the result can be readily inferred from standard RL, we provide a rigorous proof for our setting: sparse-reward, finite-horizon, episodic, and non-stationary MDPs where we maximize a GRPO-style objective. Please refer to Appendix C.2 for the full proof. ∎
@@
where $D(p_{g_{\phi}}||p_{g_{\rm{ref}}})$ should be empirical and tractable estimator of $D_{\mathrm{KL}}\!\left(p_{g_{\phi}}({\bm{x}}_{0}|{\mathbf{q}})\,\|\,p_{g_{\mathrm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})\right)$ and clipped loss has been proven to attain global optimality (Huang et al., 2024). Please note that $(g)$ in superscript of Eq. (21) is the group index.
@@ 4.1 Experimental Setting
Generation protocols and rewards. We utilize LLaDA-8B-Instruct (Nie et al., 2025) for all experiments. For logic puzzles, the model receives a prompt and a masked solution; it then denoises by unmasking one position at a time. The reward is the fraction of slots predicted correctly (e.g., in Sudoku, the proportion of blank cells filled with the correct digits). For mathematical reasoning, the model is given a problem and denoises a length-128 masked sequence. Following the LLaDA generation protocol, we partition the 128 positions into four bins of size 32 and perform any-order sampling sequentially over the bins. The reward is $1$ if the final answer matches the ground truth and 0 otherwise.
@@ 5 Conclusion
Future directions. On Sudoku, max-confidence exceeds Top-$K$, suggesting that near-optimal training can benefit from occasional low-confidence selections. Meanwhile, on GSM8K, the much larger search space makes Top-$K$ preferable for stability. In future work, we plan to develop an unmasking policy optimization algorithm that explores the full mask tokens even when the search space is large.

### v2

@@ Abstract
Masked diffusion models (MDMs) have recently emerged as a novel framework for language modeling. MDMs generate sentences by iteratively denoising masked sequences, filling in [MASK] tokens step by step. Although MDMs support any-order sampling, performance is highly sensitive to the choice of which position to unmask next. Prior work typically relies on rule-based schedules (e.g., max-confidence, max-margin), which provide ad hoc improvements. In contrast, we replace these heuristics with a learned scheduler. Specifically, we cast denoising as a KL–regularized Markov decision process (MDP) with an explicit reference policy and optimize a regularized objective that admits policy-improvement and convergence guarantees under standard assumptions. We prove that the optimized policy under this framework generates samples that more closely match the data distribution than heuristic schedules. Empirically, across four benchmarks, our learned policy consistently outperforms max-confidence: for example, on Sudoku, where unmasking order is critical, it yields a 20.1% gain over random and a 11.2% gain over max-confidence. Code is available at https://github.com/chunsanHong/UPO.
@@ 1 Introduction
Related works. Reinforcement learning (RL) has emerged as an effective tool to fine-tune large language models toward complex objectives or preferences (Ouyang et al., 2022; Guo et al., 2025; Shao et al., 2024; Rafailov et al., 2023). In MDMs, Zekri and Boullé (2025) apply policy gradient methods to a pretrained discrete diffusion model so as to maximize a task-specific reward at the last denoising step. Zhao et al. (2025) adopts the GRPO framework to improve the reasoning ability by rewarding only the final answer accuracy. Zhu et al. (2025) propose a variance reduced preference optimization (VRPO) for MDMs, which shares the concept of direct preference optimization (DPO) (Rafailov et al., 2023) in ARMs. While the aforementioned works utilize RL to post-train the MDM itself, they differ from ours, where we train the unmasking policy model. Recently, Huang et al. (2025) introduce DColT, which also trains an unmasking policy via reinforcement learning. Our approach differs in that we formulate denoising as a KL–regularized MDP with an explicit reference policy and optimize a regularized objective that admits policy-improvement and convergence guarantees under standard assumptions. Furthermore, while Huang et al. (2025) jointly applies RL to train both the unmasking policy and the underlying MDMs for maximum accuracy, our work focuses solely on optimizing the unmasking policy for frozen MDMs, thereby reducing computational cost and mitigating the risk of reward over-optimization.
@@ 2 Preliminary
Notations. Let the distribution $p_{\rm{data}}$ be the data distribution over sequences of length $L$ with vocabulary $\mathcal{V}=\{1,...,m\}$, and let the set of all possible sequences be $\mathcal{X}=\mathcal{V}^{L}$. We define $p_{L}$ as the Kronecker delta distribution that assigns all probability mass to the sentence in which every token is masked. We denote the kernel forms of $p_{\rm{data}}$ and $p_{L}$ by ${\mathbf{P}}_{\rm{data}}$ and ${\mathbf{P}}_{L}$, respectively, with entries ${\mathbf{P}}_{\rm{data}}]{{\bm{x}}}=p_{\rm{data}}({\bm{x}})$. We write ${\bm{x}}_{n}=\{x_{n}^{1},x_{n}^{2},\dots,x_{n}^{L}\}$ for the $n$-th step sequence with $L$ tokens. We denote the mask by $\mathbb{M}$ and the set of sequences with $n$ masks by $\mathcal{X}_{n}=\{{\bm{x}}|{\bm{x}}\in\{\mathcal{V}\cup\{\mathbb{M}\}\}^{L},\sum_{i=1}^{L}{\textbf{1}(x^{i}=\mathbb{M})}=n\}$, so that ${\bm{x}}_{n}\in\mathcal{X}_{n}$ has $n$ masks and we refer $a^{i}[{\bm{x}}]$ for the $i$-th mask index of $\bm{x}$}. For example, if the number of masks in $\bm{x}$} is $n$, then $a^{i}[{\bm{x}}]\in\{1,\dots,L\}$, $x^{a^{i}[{\bm{x}}]}=\mathbb{M}$ and $i\in\{1,\dots,n\}$. For brevity, we will omit $\bm{x}$} in $a^{i}[{\bm{x}}]$ for the rest of the paper. We denote MDMs as $\theta$ and their model distribution as $\pi_{\theta}$. We denote Markov transition kernel ${\mathbf{T}}:\mathcal{X}\times\mathcal{X}\to[0,1]$ where ${\mathbf{T}}({\bm{x}},{\bm{x}}^{\prime})$ is the probability of $\bm{x}$} transform into $\bm{x}$$^{\prime}$ such that $\sum_{{\bm{x}}^{\prime}\in\mathcal{X}}{\mathbf{T}}({\bm{x}},{\bm{x}}^{\prime})=1$ for all $\bm{x}$}.
@@
Now consider the MDM directly answer the questions, filling the mask of sentence: “House 1: name: [MASK], food: [MASK], House 2: name: [MASK], food: [MASK]”. Obviously, if MDM tries to answer the name of the person who lives in the first house, the problem becomes trivial. However, if it tries to fill attributes for the second house first before filling attributes of the first house, the problem becomes much harder. As the number of questions and clues grows, this kind of discrepancy would be much larger. Indeed, the zebra puzzle is also known as an NP-hard problem (Shah et al., 2024; Yato and Seta, 2003). Note that this is a highly simplified example to help understand the problem of MDM training, and refer to Appendix A for more details and Kim et al. (2025) for a complete explanation.
@@ 3 Improving Unmasking Policy Beyond a Reference Policy
3.1 Existence of optimal path Beyond $g_{\mathrm{conf}}$
@@
We test various $g$ in GSM8K (Cobbe et al., 2021), the mathematical questions benchmark, to probe whether max-confidence is merely strong or whether substantially better unmasking paths exist. We compare three schedulers: $g_{\rm conf}$, $g_{\rm rand}$, and $g_{\rm{Top-K}}$. Here, $g_{\rm conf}$ is deterministic, while $g_{\rm Top\text{-}K}$ smoothly interpolates between random and max-confidence. For controlled comparison to analyze the effect of generation ordering and following the default setting used in LLaDA (Nie et al., 2025), we conducted token sampling with argmax operator, i.e., $x_{n-1}^{a_{n}}=\operatorname*{arg\,max}_{x\in\mathcal{V}}\pi_{\theta}(x^{a_{n}}=x|\mathbf{x}_{n})$.
@@ 3.3 Theoretical analysis
Assume there exists an ideal MDM $\theta^{*}$ satisfying MDM $\mathcal{L}_{\text{MDM}}(\theta^{*})=0$ for $p_{\text{data}}$ where data is composed of prompt and oracle answer $\{\mathbf{q},\mathbf{a}\}$. Assume we select well-defined reward $r(\mathbf{q},\mathbf{a})\in\{0,1\}$ such that $p_{\text{data}}(\mathbf{q},\mathbf{a})>0\Rightarrow r(\mathbf{q},\mathbf{a})=1$ and $0<r_{g_{\text{ref}}}<1$. Now consider incomplete MDM $\theta$ such that $\mathcal{L}_{\rm{MDM}}(\theta)>0$, and $g_{\phi^{*}}$ that satisfies Eq. (15) for $\theta$. Let $g_{\mathrm{ref}}$ be a policy model such that, for every $\mathbf{q}$}, $p_{g_{\rm{ref}}}({\bm{x}}_{0}\mid{\mathbf{q}})>0$ whenever $p_{\theta^{*}}({\bm{x}}_{0}\mid{\mathbf{q}})>0$, i.e., $\operatorname{supp}(p_{\theta^{*}}(\cdot\mid{\mathbf{q}}))\subseteq\operatorname{supp}(p_{g_{\mathrm{ref}}}(\cdot\mid{\mathbf{q}}))$. Then, the following inequality holds:
$D_{\rm{KL}}(p_{\rm{\theta^{*}}}({\bm{x}}_{0}|{\mathbf{q}})||p_{g_{\phi^{*}}}({\bm{x}}_{0}|{\mathbf{q}})){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}<}D_{\rm{KL}}(p_{\rm{\theta^{*}}}({\bm{x}}_{0}|{\mathbf{q}})||p_{g_{\rm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})).$ (16)
@@
$D_{\rm{KL}}(p_{\rm{data}}({\mathbf{a}}|{\mathbf{q}})||p_{g_{\phi^{*}}}({\bm{x}}_{0}|{\mathbf{q}})){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}<}D_{\rm{KL}}(p_{\rm{data}}({\mathbf{a}}|{\mathbf{q}})||p_{g_{\rm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})).$ (17)
@@
Mroueh (2025) provides policy dynamic under training Eq. (12). Using the result given in Lemma 1 in Appendix, we can express $g_{\phi^{*}}$ with $g_{\rm{ref}}$ and thereby derive KL divergence. Refer to Appendix C.1 for detailed proof and reason why the support assumption holds. ∎
The interpretation of Theorem 2 is as follows. If we choose a binary reward $r(\mathbf{q},\mathbf{a})\in\{0,1\}$ satisfying $p_{\text{data}}(\mathbf{q},\mathbf{a})>0\Rightarrow r(\mathbf{q},\mathbf{a})=1$ and $0<r_{g_{\text{ref}}}<1$, then the theorem guarantees that the KL divergence is strictly tightened compared to the reference policy. Consequently, one may view any reward that satisfies these conditions as a “good” reward in the sense that it provably induces learning dynamics that bring the sampling distribution closer to $p_{\text{data}}$. Under the assumption that the dataset is well constructed, such binary rewards naturally correspond to properties such as correctness, the absence of reasoning errors, grammatical validity, or even combinations thereof, any of which encourage the model to move its output distribution closer to the true data distribution.
@@ 3.4 Practical realization of unmasking policy optimization
Although the result can be readily inferred from standard RL, we provide a rigorous proof for our setting: sparse-reward, finite-horizon, episodic, and non-stationary MDPs where we maximize a GRPO-style objective. The full proof is in Appendix C.2. ∎
@@
where $D(p_{g_{\phi}}||p_{g_{\rm{ref}}})$ should be empirical and tractable estimator of $D_{\mathrm{KL}}\!\left(p_{g_{\phi}}({\bm{x}}_{0}|{\mathbf{q}})\,\|\,p_{g_{\mathrm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})\right)$ and clipped loss has been proven to attain global optimality (Huang et al., 2024). Note that $(g)$ in superscript of Eq. (21) is the group index.
@@ 4.1 Experimental Setting
Generation protocols and rewards. We utilize LLaDA-8B-Instruct (Nie et al., 2025) for all experiments. For logic puzzles, the model receives a prompt and a masked solution; it then denoises by unmasking one position at a time. The reward is the fraction of slots predicted correctly (e.g., in Sudoku, the proportion of blank cells filled with the correct digits). For mathematical reasoning, the model is given a problem and denoises a length-128 masked sequence. Following the LLaDA generation protocol, we partition the 128 positions into four bins of size 32 and perform sampling by each baselines sequentially over the bins. The reward is $1$ if the final answer matches the ground truth and 0 otherwise. In addition, we incorporate the sum of log-probabilities produced by $\pi_{\theta}(\cdot)$ as an auxiliary reward signal. Detailed explanation and settings such as training hyper-parameters are provided in Appendix E. We employ the dense reward rather than the binary reward because it leads to faster convergence and higher accuracy in practice. Nonetheless, we also verify that our proposed method outperforms the max-confidence baseline under the binary reward and report these results in Appendix D.1.
@@ 5 Conclusion
Discussion.
Our learned unmasking policy improves performance across benchmarks, though the gains on GSM8K and Math500 are relatively smaller than on Sudoku. We conjecture that this stems from two differences between the domains. First, Sudoku exhibits relatively clear and consistent ordering structure—for example, certain positions are statistically more constraining—whereas such regularity is much weaker in mathematical reasoning tasks. Second, while Sudoku requires learning ordering patterns over a small digit space, GSM8K requires learning over the full vocabulary, despite having far fewer training examples. These factors likely make it more challenging to extract reliable ordering cues in GSM8K, resulting in more modest gains.
Future directions.
Our work demonstrates that GRPO-based training can yield task-specific unmasking policies that outperform max-confidence and other rule-based heuristics. As discussed above, an important next step is to develop more generalizable policies that remain effective on larger and more diverse language datasets; we explore this direction further in Appendix D.2 with additional experiments. Another promising direction is to design more stabilized policy optimization methods. On Sudoku, the full reparametrization $g_{\phi}$ outperforms $g_{\phi_{\text{Top-K}}}$, suggesting that near-optimal training benefits from occasionally selecting low-confidence positions. In contrast, the much larger search space in GSM8K favors the stability of the Top-K–restricted policy. Developing an optimization scheme that enables effective exploration of all masked positions, even in large search spaces, constitutes an important avenue for future work.
Acknowledgments
This work was supported by the National Research Foundation of Korea under Grant RS-2024-00336454, by the Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2025-02304967, AI Star Fellowship(KAIST) & RS-2024-00457882, AI Research Hub Project), and by AI Computing Infrastructure Enhancement (GPU Rental Support) User Support Program funded by the Ministry of Science and ICT (MSIT), Republic of Korea (RQT-25-120217).
Ethics Statement
We follow the ICLR Code of Ethics. Our work trains a lightweight unmasking policy while keeping the base MDM frozen, and evaluates only on public benchmarks (Sudoku, Zebra, GSM8K, Math500) that contain no personally identifiable information; no human-subject data were collected and no IRB approval was required. As our method optimizes only a lightweight scheduling policy to maximize a verifiable terminal reward and does not modify the base MDM’s parameters or training data, the safety and fairness properties are largely inherited from the underlying model. Absent defects in the base model, the incremental risk is limited to potential amplification of existing behaviors; in practice, this can be mitigated by pairing our scheduler with standard safety filters and auditing the combined system on downstream tasks.
Reproducibility statement
The training algorithm is detailed in Algorithm 2, the overall experimental setup is described in Section 4, and the hyperparameters and prompt templates used for evaluation are provided in Appendix E. Model architecture and the policy inference procedure are summarized in Appendix B. Complete proofs of all theorems and propositions stated in the main paper appear in Appendix C.
@@ Appendix A Related works
Remark. Detailed Explanation of Hardness of MDM Training.
Kim et al. (2025) show that the conditional prediction tasks induced by Any-order masked diffusion models fall into the planted constrained satisfaction problem (CSP) hard phase (Krzakala and Zdeborová, 2009), implying that no polynomial-time algorithm can achieve optimal performance in general. Consequently, MDM training loss, Eq. (1) (the reverse-process KL objective driven by Sahoo et al. (2024); Zheng et al. (2025)) cannot be driven to zero, which is consistent with their empirical finding that loss remains high on text and L&O-NAE-SAT datasets.